OpenStudy (anonymous):

using loarithms solve:the population of a city is growing at a rate proportional to its population. the population 20 years ago was 100 000 and today it is 150 000. find the population 20 years from now

OpenStudy (anonymous):

are you supposed to use \[P(t)=100,000 e^{rt}\] or an different way?

OpenStudy (anonymous):

we can solve this without logarithms if you like

OpenStudy (anonymous):

I need to apply logs

OpenStudy (amistre64):

20 years in the future i assume

OpenStudy (amistre64):

\[\frac{dP}{dt}=P\] \[\frac{dP}{P}=\ dt\] \[ln|{P}|=t\] maybe

OpenStudy (amistre64):

just as an exercise in futility .... ln|P| = t+c ln|100 000| = -20 + C C = ln(100 000)+20 ln|P| = 20 + 20 + ln(100 000)

OpenStudy (amistre64):

close, but no bananas

OpenStudy (anonymous):

we can use \[\frac{150,000}{100,000}=1.5\] to get a simple answer of \[P(t)=100,000(1.5)^{\frac{t}{20}}\] where t is years after 20 years ago

OpenStudy (anonymous):

actually we dont even need that. we know it increases by 50% every 20 years, so increase 150,000 by 50% to get the answer in one step

OpenStudy (anonymous):

the "use logarithms" thing threw me off

OpenStudy (amistre64):

me too, i got no idea how to do it with logs

OpenStudy (amistre64):

might as well be proximate power and sufficient grace to me lol ... thnx Pascal

OpenStudy (anonymous):

i think maybe you are supposed to model this as \[G(t)=P_0e^{rt}\] and you would find r using logs, but it is silly to do such an easy problem in so many steps

OpenStudy (anonymous):

should i use this formula?